# How Reliable is Language Model Micro-Benchmarking?

### v1

@@ 2 Micro-Benchmarking Preliminaries
We consider four micro-benchmark selection methods throughout this paper. Anchor Points (Vivek et al., 2024) first calculates correlations between example pairs using source model confidences and then selects the centroids of test example clusters in the resulting embedding space, in order to obtain a high correlation between model rankings on the full benchmark and the selected micro-benchmark. The tinyBenchmarks IRT method (Polo et al., 2024) instead aims to select a micro-benchmark that minimizes the error when predicting accuracy for individual source models. It does so by training an Item Response Theory (IRT) model, which results in example and source model embeddings. It then selects the closest example to centroids obtained from clustering these embeddings. [^1] Polo et al. (2024) propose method variants that also incorporate IRT model predictions. We focus on their core method, as our initial experiments showed similar results for variants. Stratified random sampling (Fogliato et al., 2024b) randomly selects examples from clusters obtained based on model confidence on examples. We also consider a diversity-based method that samples examples spread evenly in the space of source model correlations used by Anchor Points, enabled by a sampler that can select negatively-dependent samples (Bardenet et al., 2024). Appendix A gives detailed descriptions of these methods.
@@ 3 MDAD: A Meta-Evaluation for Micro-Benchmark Reliability
The agreement function can be summarized using a single value that we call the Minimum Detectable Ability Difference (MDAD): what is the lowest performance difference on the full benchmark at which pairwise model rankings on a micro-benchmark are consistent with those on the full benchmark? Following conventions from statistical power analysis (Card et al., 2020; Cohen, 1962), we consider a judgment consistent if it is correct at least 80% of the time: [^2] Appendix B shows qualitatively similar results for different thresholds.
@@
In practice, we report the centroid of the bucket corresponding to the lowest performance difference. MDAD captures how well a micro-benchmark preserves pairwise model rankings, unlike mean estimation error, which only considers single models, or rank correlation, which considers all model rankings in the aggregate. Lower MDAD is better because then even small performance differences across model pairs can be reliably measured under $D_{\text{micro}}$. If two models’ performances differ by less than a micro-benchmark’s MDAD, then that micro-benchmark is not likely to be able to consistently distinguish them. For example, suppose a micro-benchmark at a given size results in an MDAD of 10.
Figure 2: Agreement and MDAD measures on MMLU-Pro for uniform random sampling and Anchor Points with 300 source models. The three left panels show the probability that a pairwise ranking of models on a micro-benchmark agrees with the full benchmark’s ranking, as a function of the accuracy difference between those models on the full benchmark. The rightmost panel summarizes all these agreement curves by showing the minimum detectable accuracy difference between models at each micro-benchmark size, i.e. the accuracy difference at which each curve in the first three panels crosses the 0.8 probability of agreement threshold. Points A-F show how each agreement curve is summarized by MDAD: each point marks where an agreement curve surpasses 0.8 probability. For MDAD, lower values are better. Error bars show 95% bootstrap confidence intervals over 50 trials.
Then at least 80% of the time, that micro-benchmark can correctly rank only model pairs that differ on the full benchmark by at least 10 performance points. If we instantiate agreement (Equation 4) using accuracy as the performance measure, Equation 5 then measures the Minimum Detectable Accuracy Difference. Figure 2 provides an illustrative example of the relationship between agreement at various accuracy differences and MDAD, which summarizes the agreement curve. For the rest of the paper, MDAD will refer to this accuracy-based instantiation.
@@ 4 Experimental Design
For all experiments, we simulate different draws from a benchmark by splitting each benchmark in half (each subtask is divided in half as well): the train half is used to select the micro-benchmark, and the held-out half is used to measure generalization. We also uniformly at random partition a set of models into source models for selecting micro-benchmarks and a set of target models whose accuracy we are predicting, as in Vivek et al. (2024) and Polo et al. (2024). We account for variance in micro-benchmark construction by averaging over random samples of datasets and source models, following Card et al. (2020) and Perlitz et al. (2024). Most of our experiments select a micro-benchmark from an entire benchmark (as in Polo et al. (2024)), though in §5.4 we also evaluate selecting micro-benchmarks per subtask (as in Vivek et al. (2024)) and report averages across subtasks. We compute meta-evaluation measures with the same data split used for micro-benchmark selection for most experiments, following prior work. In §5.4, we do so using the held-out set. We use 50 trials, each with a partition of data points and models. Details are in Appendix C. Appendix D analyzes MDAD estimates for up to 100 trials; we find MDADs have stabilized by 50 trials.
@@
Following prior work, we consider micro-benchmarking methods specifically designed for classification tasks. We report accuracy as a percent from 0 to 100 and measure agreement using accuracy difference buckets at a resolution of 0.5 points of accuracy, i.e. $\mathcal{B}=\{[0,0.25),[0.25,0.75),[0.75,1.25),\ldots\}$, reporting MDAD as the bucket centroid. [^3] Appendix E shows that bucket resolutions of 0.25, 0.5, and 1 all yield similar MDAD values.
@@
We evaluate constructing micro-benchmarks at various small sizes by selecting $n\in\{10,25,50,100,250,500,1000\}$ examples. For GPQA, a smaller benchmark, we select $n\in\{10,25,50,100,200\}$ examples. In Appendix F, we also evaluate constructing micro-benchmarks at various proportions of the original benchmark by selecting $\{2\%,4\%,8\%,16\%,32\%,40\%\}$ of examples, finding qualitatively similar results.
@@
Our experiments do not attempt to find the “best” examples for evaluation—rather, they are designed to assess the reliability of existing micro-benchmarking methods. We also specifically seek to understand the conditions under which random sampling is an effective alternative to existing micro-benchmarking. For this reason, we do not release specific subsets of benchmarks like tinyMMLU (Polo et al., 2024) or Flash-HELM (Perlitz et al., 2024).
Figure 3: Comparing six micro-benchmarking approaches on two benchmarks. $y$-axis shows agreement (Equation 4), the probability that a micro-benchmark agrees with the full benchmark when comparing two models, as a function of how much those models differ on the full benchmark ($x$-axis). The rightmost column summarizes agreement curves using MDAD (Equation 5). At small micro-benchmark sizes, all methods struggle to compare models that differ by fewer than 4 points of accuracy on the full benchmark. Anchor Points does best, followed by tinyBenchmarks. Error bars show 95% bootstrap confidence intervals over 50 trials. Figure 9 (Appendix F) shows all benchmarks.
@@ 5 Results
We consider four micro-benchmark selection methods and two random sampling baselines across the four benchmark suites, using agreement (Eq. 4) and MDAD (Eq. 5) from §3. MDAD reveals limitations of all the micro-benchmarking approaches we consider in the extreme dataset reduction regime and provides a finer-grained analysis than existing meta-evaluation measures (§5.1, §5.3). We show that random sampling is competitive with other micro-benchmark selection if sampling at least 250 examples (§5.2). §5.4 analyzes micro-benchmark generalization to new draws of the task. The complete results per benchmark for all parameter settings are in Appendix F. Appendix G shows that our analysis also holds for a micro-benchmarking method that selects whole subtasks.
@@
Figure 3 examines how agreement varies with micro-benchmark size across methods and benchmarks. Each agreement curve is summarized by an MDAD value (Figure 3, rightmost column); lower MDAD values correspond to higher reliability. The leftmost column of panels shows that BBH micro-benchmarks of size 10 cannot reliably rank model pairs unless they differ by almost 15 points of accuracy! The only exception to this is Anchor Points, which has higher agreement at smaller micro-benchmark sizes. As more examples are selected, model pairwise rankings on micro-benchmarks are likely to be more predictive of those on the full eval set. Just as agreement steadily increases as more examples are selected, MDAD steadily decreases for all methods as more examples are selected. When 10 examples are selected from BBH, Anchor Points achieves an MDAD of 6. If a model pair differs by more than 6 points on the full BBH then this micro-benchmark is likely to correctly rank these models. If a model pair has an accuracy difference on the full BBH of less than the MDAD, e.g. 2 points of accuracy, then this micro-benchmark is unlikely to correctly rank these models. Overall, no micro-benchmarking method can consistently distinguish models that differ on the full benchmark by fewer than 3 points of accuracy on MMLU, 3.5 points of accuracy on MMLU-Pro, 6 points of accuracy on BBH, or 6.5 points of accuracy on GPQA when selecting 10 examples.
@@ 5.1 MDAD affords finer-grained analysis than existing measures
Figure 4 compares MDAD to mean estimation error and Kendall’s tau rank correlation using the experimental design from §4. MDAD provides a complementary—but not contradictory—perspective to these existing measures. We also report the correlation between MDAD and other meta evaluation measures, using Kendall’s tau rank correlation. [^4] Note here that we are using Kendall’s tau rank correlation in a new way, to compare MDAD to existing meta-evaluation measures. One of these measures is itself the Kendall’s tau rank correlation between micro-benchmark and full benchmark model rankings that we have been considering throughout. Each setting with a combination of benchmark, micro-benchmark size, and micro-benchmarking method yields a set of meta-evaluation measure values. We calculate the correlation between the rankings of settings induced by those values. Across all settings, MDAD has a $\tau=0.701\;(p<0.05)$ Kendall’s tau rank correlation with mean estimation error. MDAD also has a $\tau=-0.787\;(p<0.05)$ correlation with the Kendall’s tau rank correlation evaluation measure. Despite the high correlation and similar trends at a high level, e.g. Anchor Points stands out on both MDAD and rank correlation, there are still fine-grained differences between the measures. MDAD does not map one-to-one to other measures.
@@
Even when micro-benchmarks have similar rank correlations in the aggregate, MDAD can identify fine-grained differences between methods. For example, tinyBenchmarks and uniform random sampling both have identical rank correlations when selecting 10 examples from MMLU-Pro (Points A and B, Fig. 4). But uniform random sampling has a much higher MDAD than tinyBenchmarks in this setting (C and D, Fig. 4). Comparing across datasets shows that different rank correlations can map to the same MDAD value. When selecting 10 examples, Anchor Points has a much higher rank correlation of 0.73 on BBH but a rank correlation of 0.43 on GPQA (E and F, Fig. 4). In both cases, MDAD is 6 (G and H, Fig. 4). Even though the rank correlations are very different, both micro-benchmarks afford consistently accurate model comparisons when the models differ by at least 6 points of accuracy. Considering model comparisons only in the aggregate, as rank correlation does, obscures this finer-grained analysis.
@@
Mean estimation error is defined for a single model; it does not take into account whether a micro-benchmark consistently overestimates or consistently underestimates model accuracy for different models. For a simple illustrative example, if mean estimation error is 5 points, then the micro-benchmark’s accuracy for a single model will be off by 5 points on average. If when comparing two models, the micro-benchmark overestimates both of their accuracies by 5 points, the micro-benchmark can still yield the correct pairwise model ranking. MDAD accounts for this by directly measuring whether pairs of models are correctly ranked. For example, when 100 examples are selected from MMLU-Pro, Anchor Points has a higher estimation error than random sampling but a lower MDAD.
@@ 5.2 Random sampling is competitive with other methods
If we were to only use Kendall’s tau rank correlation, we would have trouble differentiating most methods from random sampling when selecting between 10 and 100 examples. At these extreme dataset reductions, methods like tinyBenchmarks and stratified sampling by confidence have similar Kendall’s tau rank correlations to random sampling, indicating that they all rank models equally well in the aggregate. But these methods have lower MDADs than random sampling, showing they can more reliably distinguish models that differ by fewer points of accuracy.

### v2

@@ 2 Micro-Benchmarking Preliminaries
We consider four micro-benchmark selection methods throughout this paper. Anchor Points (Vivek et al., 2024) first calculates correlations between example pairs using source model confidences and then selects the centroids of test example clusters in the resulting embedding space, in order to obtain a high correlation between model rankings on the full benchmark and the selected micro-benchmark. The tinyBenchmarks IRT method (Polo et al., 2024) instead aims to select a micro-benchmark that minimizes the error when predicting accuracy for individual source models. It does so by training an Item Response Theory (IRT) model, which results in example and source model embeddings. It then selects the closest example to centroids obtained from clustering these embeddings. [^1] Polo et al. (2024) propose method variants that also incorporate IRT model predictions. We focus on their core method, as our initial experiments showed similar results for variants. Stratified random sampling (Fogliato et al., 2024b) randomly selects examples from clusters obtained based on model confidence on examples. We also consider a diversity-based method that samples examples spread evenly in the space of source model correlations used by Anchor Points, enabled by a sampler that can select negatively-dependent samples (Bardenet et al., 2024). Appendix B gives detailed descriptions of these methods.
@@ 3 MDAD: A Meta-Evaluation for Micro-Benchmark Reliability
The agreement function can be summarized using a single value that we call the Minimum Detectable Ability Difference (MDAD): what is the lowest performance difference on the full benchmark at which pairwise model rankings on a micro-benchmark are consistent with those on the full benchmark? Our goal here is to adapt the idea from statistical power analysis of estimating the minimum difference in model performance that can be consistently detected by a dataset (Card et al., 2020; Cohen, 1962). Following conventions from statistical power analysis, we consider a judgment consistent if it is correct at least 80% of the time: [^2] Appendix C shows qualitatively similar results for different thresholds.
@@
In practice, we report the centroid of the bucket corresponding to the lowest performance difference. MDAD captures how well a micro-benchmark preserves pairwise model rankings, unlike mean estimation error, which only considers single models, or rank correlation, which considers all model
Figure 2: Agreement and MDAD measures on MMLU-Pro for uniform random sampling and Anchor Points with 300 source models. The three left panels show the probability that a pairwise ranking of models on a micro-benchmark agrees with the full benchmark’s ranking, as a function of the accuracy difference between those models on the full benchmark. The rightmost panel summarizes all these agreement curves by showing the minimum detectable accuracy difference between models at each micro-benchmark size, i.e. the accuracy difference at which each curve in the first three panels crosses the 0.8 probability of agreement threshold. Points A-F show how each agreement curve is summarized by MDAD: each point marks the minimum difference in accuracy where an agreement curve surpasses a 0.8 probability of agreement. For MDAD, lower values are better. Error bars represent 95% bootstrap confidence intervals over 50 trials.
rankings in the aggregate. Lower MDAD is better because then even small performance differences across model pairs can be reliably measured under $D_{\text{micro}}$. If two models’ performances differ by less than a micro-benchmark’s MDAD, then that micro-benchmark is not likely to be able to consistently distinguish them. For example, suppose a micro-benchmark at a given size results in an MDAD of 10. Then at least 80% of the time, that micro-benchmark can correctly rank only model pairs that differ on the full benchmark by at least 10 performance points. If we instantiate agreement (Equation 4) using accuracy as the performance measure, Equation 5 then measures the Minimum Detectable Accuracy Difference. Figure 2 provides an illustrative example of the relationship between agreement at various accuracy differences and MDAD, which summarizes the agreement curve. For the rest of the paper, MDAD will refer to this accuracy-based instantiation. Table 1 gives a conceptual overview of how MDAD differs from existing meta-evaluation measures.
Table 1: Summary of differences between MDAD and existing meta-evaluation measures.
Meta-evaluation measure What does it measure? Unit of comparison Aggregation
Mean estimation error Raw performance Individual model Across all models
Kendall’s tau rank correlation Model rankings Model pair Across all model pairs
MDAD Model rankings Model pair Model pairs split by performance difference
@@ 4 Experimental Design
For all experiments, we simulate different draws from a benchmark by splitting each benchmark in half (each subtask is divided in half as well): the train half is used to select the micro-benchmark, and the held-out half is used to measure generalization. We also uniformly at random partition a set of models into source models for selecting micro-benchmarks and a set of target models whose accuracy we are predicting, as in Vivek et al. (2024) and Polo et al. (2024). We account for variance in micro-benchmark construction by averaging over random samples of datasets and source models, following Card et al. (2020) and Perlitz et al. (2024). Most of our experiments select a micro-benchmark from an entire benchmark (as in Polo et al. (2024)), though in §5.4 we also evaluate selecting micro-benchmarks per subtask (as in Vivek et al. (2024)) and report averages across subtasks. We compute meta-evaluation measures with the same data split used for micro-benchmark selection for most experiments, following prior work. In §5.4, we do so using the held-out set. We use 50 trials, each with a partition of a) data points into a set for selection and a set for measuring generalization and b) models into source and target sets. Details are in Appendix D. Appendix E analyzes MDAD estimates for up to 100 trials; we find MDADs have stabilized by 50 trials.
@@
Following prior work, we consider micro-benchmarking methods specifically designed for classification tasks. We report accuracy as a percent from 0 to 100 and measure agreement using accuracy difference buckets at a resolution of 0.5 points of accuracy, i.e. $\mathcal{B}=\{[0,0.25),[0.25,0.75),[0.75,1.25),\ldots\}$, reporting MDAD as the bucket centroid. [^3] Appendix F shows that bucket resolutions of 0.25, 0.5, and 1 all yield similar MDAD values. In our experiments, MDAD takes on average 2.40 seconds to compute (with a standard deviation of 0.24 s).
@@
We evaluate constructing micro-benchmarks at various small sizes by selecting $k\in\{10,25,50,100,250,500,1000\}$ examples. For GPQA, a smaller benchmark, we select $k\in\{10,25,50,100,200\}$ examples. In Appendix G, we also evaluate constructing micro-benchmarks at various proportions of the original benchmark by selecting $\{2\%,4\%,8\%,16\%,32\%,40\%\}$ of examples, finding qualitatively similar results.
@@
Our experiments do not attempt to find the “best” examples for evaluation—rather, they are designed to assess the reliability of existing micro-benchmarks. We also specifically seek to understand the conditions under which random sampling is an effective alternative to existing micro-benchmarking. For this reason, we do not release specific subsets of benchmarks like tinyMMLU (Polo et al., 2024) or Flash-HELM (Perlitz et al., 2024).
Figure 3: Comparing six micro-benchmarking approaches on two benchmarks. $y$-axis shows agreement (Equation 4), the probability that a micro-benchmark agrees with the full benchmark when comparing two models, as a function of how much those models differ on the full benchmark ($x$-axis). The rightmost column summarizes agreement curves using MDAD (Equation 5). For small micro-benchmarks, all methods struggle to compare models that differ by fewer than 4 points of accuracy on the full benchmark. Anchor Points does best, followed by tinyBenchmarks. Error bars show 95% bootstrap confidence intervals over 50 trials. Figure 9 (Appendix G) shows all benchmarks.
@@ 5 Results
We consider four micro-benchmark selection methods and two random sampling baselines across the four benchmark suites, using agreement (Eq. 4) and MDAD (Eq. 5) from §3. MDAD reveals limitations of all the micro-benchmarking approaches we consider in the extreme dataset reduction regime and provides a finer-grained analysis than existing meta-evaluation measures (§5.1, §5.3). We show that random sampling is competitive with other micro-benchmark selection if sampling at least 250 examples (§5.2). §5.4 analyzes micro-benchmark generalization to new draws of the task. The complete results per benchmark for all parameter settings are in Appendix G. Appendix H shows that our analysis also holds for a micro-benchmarking method that selects whole subtasks.
Larger micro-benchmarks afford lower MDAD.
Figure 3 examines how agreement varies with micro-benchmark size across methods and benchmarks. Each agreement curve is summarized by an MDAD value (Figure 3, rightmost column); lower MDAD values correspond to higher reliability. As a first case study, consider BBH micro-benchmarks (Figure 3, bottom row). As more examples are selected, model pairwise rankings on micro-benchmarks are likely to be more predictive of those on the full eval set. For example, the agreement curves in the 100-example panel have shifted to the left of their positions in the 10-example panel. Just as agreement steadily increases as more examples are selected, MDAD steadily decreases for all methods as more examples are selected. As the number of examples increases from 10 to 100 to 1000, MDAD for tinyBenchmarks drops from 16 to 4 to 1.
@@
The leftmost column of Figure 3 shows that BBH micro-benchmarks of size 10 cannot reliably rank model pairs unless they differ by almost 15 points of accuracy! The only exception to this is Anchor Points, which has higher agreement at smaller micro-benchmark sizes. When 10 examples are selected from BBH, Anchor Points achieves an MDAD of 6. If a model pair differs by more than 6 points on the full BBH, then this micro-benchmark is likely to correctly rank these models. If a model pair has an accuracy difference on the full BBH of less than the MDAD, e.g. 2 points of accuracy, then this micro-benchmark is unlikely to correctly rank these models. Overall all micro-benchmarking methods are limited when selecting only 10 examples. No method can consistently distinguish models that differ on the full benchmark by fewer than 3 points of accuracy on MMLU, 3.5 points of accuracy on MMLU-Pro, 6 points of accuracy on BBH, or 6.5 points of accuracy on GPQA.
Anchor Points has the lowest MDAD at the smallest sizes but stagnates.
The rightmost column of Figure 3 shows that Anchor Points has lower MDAD than other methods when selecting extremely few examples. We suspect this is because its method of selecting examples based on correlations in source model confidence is more closely aligned with what MDAD measures. On the other hand, Anchor Points has the highest MDADs when selecting 1,000 examples. We hypothesize this is due to very imbalanced cluster sizes in the underlying clustering that it performs. Anchor Points works by first clustering all of the full benchmark’s examples using $k$-medoids and then selecting one example from each cluster. When selecting 10 examples from MMLU-Pro, cluster sizes are relatively even. But when selecting 1,000 examples, there is an extreme size imbalance between the 1,000 clusters of the benchmark examples: 47% of clusters are singletons, and the largest 10% of clusters together contain half of all the examples. Contrast this with tinyBenchmarks, which also performs clustering but uses $k$-means rather than $k$-medoids and uses a different embedding space. In the same setting of selecting 1,000 examples from MMLU with tinyBenchmarks, only 5% of clusters are singletons, and the largest 10% of clusters only contain 21% of examples. Each selected example from the large clusters in Anchor Points must stand in for many more data points.
@@ 5.1 MDAD affords finer-grained analysis than existing measures
Figure 4 compares MDAD to mean estimation error and Kendall’s tau rank correlation using the experimental design from §4. MDAD provides a complementary—but not contradictory—perspective to these existing measures. We also report the correlation between MDAD and other meta-evaluation measures, using Kendall’s tau rank correlation. [^4] Note here that we are using Kendall’s tau rank correlation in a new way, to compare MDAD to existing meta-evaluation measures. One of these measures is itself the Kendall’s tau rank correlation between micro-benchmark and full benchmark model rankings that we have been considering throughout. Each setting with a combination of benchmark, micro-benchmark size, and micro-benchmarking method yields a set of meta-evaluation measure values. We calculate the correlation between the rankings of settings induced by those values. Across all settings, MDAD has a $\tau=0.701\;(p<0.05)$ Kendall’s tau rank correlation with mean estimation error. MDAD also has a $\tau=-0.787\;(p<0.05)$ correlation with the Kendall’s tau rank correlation evaluation measure. Despite the high correlation and similar trends at a high level, e.g. Anchor Points stands out on both MDAD and rank correlation, there are still fine-grained differences between the measures. MDAD does not map one-to-one to other measures.
@@
Even when micro-benchmarks have similar rank correlations in the aggregate, MDAD can identify fine-grained differences between methods. For example, tinyBenchmarks and uniform random sampling both have identical rank correlations when selecting 10 examples from MMLU-Pro (Points A and B, Fig. 4). But uniform random sampling has a much higher MDAD than tinyBenchmarks in this setting (C and D, Fig. 4). Comparing across datasets shows that different rank correlations can map to the same MDAD value. When selecting 10 examples, Anchor Points has a much higher rank correlation of 0.73 on BBH but a rank correlation of 0.43 on GPQA (E and F, Fig. 4). In both cases, MDAD is 6 (G and H, Fig. 4). Even though the rank correlations are very different, both micro-benchmarks afford consistently accurate model comparisons when the models differ by at least 6 points of accuracy. Considering model comparisons only in the aggregate, as rank correlation does, obscures this finer-grained analysis. Additionally, MDAD values have a concrete interpretation—namely, the minimum model performance difference that an evaluation dataset can distinguish at least 80% of the time—while Kendall’s tau values are harder to interpret on their own.
@@
Mean estimation error is defined for a single model; it does not take into account whether a micro-benchmark consistently overestimates or consistently underestimates model accuracy for different models. For a simple illustrative example, if mean estimation error is 5 points, then the micro-benchmark’s accuracy for a single model will be off by 5 points on average. If when comparing two models, the micro-benchmark overestimates both of their accuracies by 5 points, the micro-benchmark can still yield the correct pairwise model ranking. MDAD accounts for this by directly measuring whether pairs of models are correctly ranked. For example, when 100 examples are selected from MMLU-Pro, Anchor Points has a higher estimation error than random sampling but a lower MDAD. Mean estimation error does not directly capture when models will be ranked correctly, but MDAD does.
@@ 5.2 Random sampling is competitive with other methods
Case study: MDAD better distinguishes micro-benchmarks from random at $\leq$100 examples.
If we were to only use Kendall’s tau rank correlation, we would have trouble differentiating most methods from random sampling when selecting between 10 and 100 examples. Consider performance on MMLU for $\leq$100 examples (Figure 4, left column). At these extreme dataset reductions, methods like tinyBenchmarks and stratified sampling by confidence have similar Kendall’s tau rank correlations to random sampling, indicating that they all rank models equally well in the aggregate. But these methods have lower MDADs than random sampling, showing they can more reliably distinguish models that differ by fewer points of accuracy. This observation also holds for the other benchmarks.